# Towards deep physics-informed Kolmogorov–Arnold networks

### v1

@@ 3.2 Small-Scale Benchmarks
To evaluate the effectiveness of the proposed initialization scheme, we compare it against the prevailing practice for Chebyshev-based KANs. In this default approach, the basis-function coefficients $w_{jim}$ are drawn from a normal distribution with zero mean and variance $\big[d_{\text{I}}(D+1)\big]^{-1}$ [31, 42]. To this end, we conduct experiments on two small-scale benchmarks: (i) function fitting, and (ii) forward PDE problems.
@@ 3.2.1 Function Fitting
We first assess initialization performance on five function-fitting tasks of increasing dimensionality: (i) a one-dimensional oscillatory function, $f_{1}(x)$, (ii) the two-dimensional product function, $f_{2}(x_{1},x_{2})$, (iii) a more challenging two-dimensional function borrowed from [44], $f_{3}(x_{1},x_{2})$, (iv) the three-dimensional Hartmann function, $f_{4}(x_{1},x_{2},x_{3})$, which is a common benchmark in function approximation, and (v) the five-dimensional Sobol g-function, $f_{5}(x_{1},\dots,x_{5})$, widely used in global sensitivity analysis. The analytic definitions of all functions are provided in C.
For each function, we generated $4\cdot 10^{3}$ input–output samples uniformly over the domain $1,1]^{d}$, with $d$ being the dimensionality of the function. Chebyshev-based KANs with polynomial order $D=8$ were trained under both the default initialization and the proposed Glorot-like scheme. In all cases, we modulated the gain factor to account for the standard deviation of inputs sampled from $\mathcal{U}(-1,1)$, a convention we adopt throughout all subsequent experiments. Networks were trained for $2\cdot 10^{3}$ iterations using the Adam [62] optimizer with a constant learning rate of $10^{-3}$ in full-batch mode. Architectures of varying width (hidden layer dimensions of 2, 4, 8, 16, 32, 64) and depth (2 to 5 hidden layers) were considered. Training minimized the $L^{2}$ loss, and final performance was evaluated in terms of the relative $L^{2}$ error with respect to the reference solution (see B for further details). To ensure statistical significance, each configuration was repeated with five random seeds. For the final relative $L^{2}$ error evaluation, we used 1,000 uniformly spaced points in [1,1][-1,1] for the one-dimensional task, a $200\times 200$ grid for the two-dimensional tasks, a $30^{3}$ grid for the three-dimensional Hartmann function, and a $10^{5}$ grid for the five-dimensional Sobol g-function.
The results of these experiments are summarized in Figure 1. Each heatmap corresponds to one benchmark function, with hidden layer dimension on the horizontal axis and number of hidden layers on the vertical axis. The color scale quantifies the relative improvement achieved by the proposed initialization over the default scheme, computed as
@@
Figure 2: Loss throughout training for two representative architectures (top row: three 4-dimensional hidden layers; bottom row: five 16-dimensional hidden layers) across the five benchmark functions. Each subplot shows the mean training loss over five independent runs (solid lines) together with the standard error (shaded area). The final column, corresponding to $f_{5}$, is shown without logarithmic scaling on the $y$-axis, since the loss did not exhibit significant improvement during training.
Apart from the final error metrics, it is also informative to examine the training loss evolution, in order to assess whether the proposed initialization leads to a more effective optimization of the loss function. To this end, we considered two representative architectures: a smaller network with width $4$ and depth $3$, and a larger network with width $16$ and depth $5$. Figure 2 depicts the training loss curves for each of the five benchmark functions under both initialization schemes. Solid lines indicate the mean loss across five independent runs, while the shaded regions correspond to the standard error. The proposed initialization consistently accelerates convergence and achieves substantially lower training losses. For the oscillatory function ($f_{1}$) and the two-dimensional cases ($f_{2}$ and $f_{3}$), the difference spans more than two orders of magnitude, with low variability across seeds. Similar improvements are observed for the Hartmann function ($f_{4}$), where the default initialization stalls at higher loss values compared to the proposed scheme. For the Sobol $g$-function ($f_{5}$), both schemes exhibit nearly identical behavior, in line with the earlier observation that this benchmark is less sensitive to initialization. Importantly, these trends are visible in both architectures, demonstrating robustness across different model scales.
@@ 3.2.2 Forward PDE Problems
We next assess the proposed initialization on forward PDE benchmarks using cPIKANs. Specifically, we consider Burgers’ equation as well as the Helmholtz equation with $a_{1}=1$, $a_{2}=4$ (see D for details), following the PIML framework introduced in Section 2, without incorporating any additional adaptive techniques.
For each PDE, $N_{\text{pde}}=2^{12}$ collocation points are used to enforce the differential operator. In the Burgers’ case, $N_{\text{bc}}=2^{6}$ points are used for each of the two boundary conditions, together with $N_{\text{ic}}=2^{6}$ points for the initial condition. For Helmholtz, $N_{\text{bc}}=2^{6}$ points are used for each of the four boundary conditions. Collocation points are sampled once from a uniform grid and remain fixed throughout training. The trained Chebyshev-based KANs use polynomial order $D=8$ and are initialized with either the default or the proposed scheme. Training is performed for $5\cdot 10^{3}$ iterations using the Adam optimizer with a constant learning rate of $10^{-3}$ in full-batch mode. Architectures of varying width (2, 4, 8, 16, 32, 64 units per hidden layer) and depth (2 to 5 hidden layers) are considered. Models are evaluated in terms of the relative $L^{2}$ error with respect to the reference solution, and each experiment is repeated with five different random seeds.
Figure 3: Comparison of default and proposed initialization schemes on Burgers’ (top row) and Helmholtz (bottom row) equations. Left column: heatmaps of relative improvement in final $L^{2}$ error. Middle/Right column: training-loss curves per initialization scheme for a representative architecture of depth = 3/5 and width = 4/16, respectively. Shaded regions denote standard error across five runs.
The results of these experiments are summarized in Figure 3, which combines heatmaps of final relative $L^{2}$ error improvements with representative training-loss curves. As in the function-fitting benchmarks, the proposed initialization consistently outperforms the default scheme in terms of final error, with the exception of only two small architectures where the default initialization yields a marginal advantage. The overall gains are somewhat less pronounced than in Section 3.2.1, which can be attributed to the relatively low number of training iterations and the absence of adaptive training techniques. Nevertheless, the training-loss curves reveal a striking effect of initialization. In particular, for the Helmholtz equation, where it is well known that fast convergence usually requires adaptive weighting of PDE and boundary condition terms (see, e.g., [35]), the proposed initialization achieves up to eight orders of magnitude lower training loss compared to the default scheme. This difference corresponds not only to better convergence but, in practice, to the distinction between KANs that diverge under the default initialization and KANs that successfully approximate the true solution.
@@ 3.3 Training Divergence with Increasing Depth
The benchmarks considered so far involved relatively shallow networks, consistent with most current KAN applications, where architectures are typically limited to a small number of hidden layers. Within this setting, the proposed initialization was shown to improve accuracy and training stability across both function-fitting and forward PDE tasks. A natural question, however, is whether these gains extend to deeper architectures. To this end, we revisit Burgers’ equation and additionally consider the Allen–Cahn equation (see D for details), training cPIKANs of increasing depth to examine how network depth influences training stability under both initialization schemes.
In contrast to the previous benchmarks, here we adopt all of the adaptive training methods described in Section 2.3, since both PDEs are trained for a large number of iterations and the Allen–Cahn equation in particular fails to yield accurate results without adaptive strategies [25, 26]. Throughout training, adaptive collocation-point resampling is performed, with hyperparameters $\delta=1$ and $C=1$, which is a standard choice in the literature [51]. The pool of collocation points to enforce the PDE is constructed using $N_{\text{pool}}=400\times 400$ uniformly distributed samples over the spatiotemporal domain $0,1]\times[-1,1]$. Every $2\cdot 10^{3}$ iterations, $N_{\text{pde}}=2^{12}$ points are resampled for training. Alongside this, we apply the RBA method with hyperparameters $\gamma=0.999$ and $\eta=0.01$ as in [26], so that resampling follows the probability density function defined in Eq. (25). For the initial condition, a fixed set of $N_{\text{ic}}=2^{6}$ points is used, while RBA weighting is still applied. No collocation points are used to enforce boundary conditions, which are instead incorporated directly into the network architecture following the exact boundary enforcement strategy of [63].
In addition, every $10^{3}$ iterations we apply learning-rate annealing according to Eqs. (19)-(20), with decay parameter $a=0.9$. Causal training is also employed by partitioning the temporal domain into $M=32$ segments and using $\epsilon=1.0$ [23]. In all experiments, Chebyshev-based KANs with polynomial order $D=5$ are considered, initialized under both the default and the proposed scheme. Models are trained for $10^{5}$ iterations using the Adam optimizer with an initial warm-up phase of $10^{3}$ iterations, reaching a learning rate of $10^{-3}$, followed by exponential decay with a factor of $0.9$ every $2\cdot 10^{3}$ iterations. Architectures with $\{2,4,6,8,10,12\}$ hidden layers (depths) and hidden layer sizes (widths) $\{8,16,32\}$ are considered. Performance is evaluated in terms of the relative $L^{2}$ error with respect to the reference solution, and each experiment is repeated five times with different random seeds.
The results of these experiments are summarized in Figure 4, which presents the relative $L^{2}$ error of each network configuration under both initialization schemes for the two studied PDEs. Each column corresponds to a specific network width, with the horizontal axis of each subplot representing the network depth. Solid lines denote the mean relative $L^{2}$ error over the five independent runs, and the shaded regions indicate the corresponding standard error. The results highlight the consistent advantage of the proposed initialization across all architectures and both PDEs, often yielding relative $L^{2}$ errors lower by several orders of magnitude compared to the default scheme. For the Burgers’ equation, the effect is particularly striking: beyond a depth of four hidden layers, models initialized with the default scheme exhibit complete divergence, with relative errors of $\mathcal{O}(1)$, whereas the proposed initialization maintains stable training and achieves errors as low as $\mathcal{O}(10^{-3})$$\mathcal{O}(10^{-4})$. Although some degradation in accuracy is observed at larger depths for the widest networks (width $32$), no divergence occurs, and the relative errors remain well below those of the default scheme. In contrast, for the Allen–Cahn equation, training instability persists for deeper networks under both initialization schemes. While the proposed initialization consistently improves performance compared to the default, the relative error grows rapidly with depth and reaches $\mathcal{O}(1)$ for the deepest configurations. The lowest errors are typically achieved at shallow depths (i.e., depth $2$ for width $32$ and depth $4$ for the other two cases), after which performance degrades, albeit not monotonically for all cases. Overall, while the proposed scheme substantially improves training stability and accuracy, it does not, by itself, guarantee stable convergence in all cases as network depth increases.
@@
Figure 4: Relative $L^{2}$ error across increasing network depths for Burgers’ (top row) and Allen–Cahn (bottom row) equations, under both default and proposed initialization schemes. Each column corresponds to a different network width (8, 16, 32). Solid lines show mean values over five random seeds, while shaded areas represent standard error.
@@ 4 Residual-Gated Adaptive KANs
While the proposed Glorot-like initialization improved training stability and accuracy across all benchmarks, its effectiveness in mitigating depth-related issues proved case-dependent. In particular, for Burgers’ equation, cPIKANs initialized with the proposed scheme remained stable and convergent even at larger depths, whereas for the Allen–Cahn equation similar architectures exhibited divergence beyond a certain number of hidden layers. A similar phenomenon is observed in PINNs: when networks are initialized following the Glorot scheme, training tends to diverge as depth increases, regardless of the specific activation function among those commonly used in practice [23]. However, before attempting to transfer to cPIKANs the remedies that have been proposed for this behavior in PINNs, it is first necessary to determine whether the underlying mechanisms are indeed analogous.
@@ 4.1 KAN Derivatives at Initialization
In the case of MLP-based networks – the backbone of PINNs – [23] demonstrated that training divergence with increasing depth originates from the behavior of the network’s derivatives at initialization. To illustrate this, they considered a simplified MLP with scalar input and output, employing the hyperbolic tangent activation $\sigma(x)=\tanh(x)$, and focused on the first-order derivative with respect to the input. At initialization, the network operates in a near-linear regime where $\sigma(x)\approx x$, leading to the following approximation for the first-order derivative of Eq. (10) with respect to the input coordinate $x$:
@@
For Chebyshev-based KANs, where the basis functions are given by Eq. (16), it can be shown (see A.2 for the detailed derivation) that, in the linear regime,
@@ 4.2 Proposed Architecture
To address these instabilities in PINNs, [23] introduced the PirateNet architecture (see E for details). PirateNets incorporate several architectural components known to improve accuracy, such as random Fourier feature (RFF) embeddings [64] and a physics-informed initialization of the final network layer. However, the key idea for resolving the depth-related issue is the introduction of an adaptive skip connection. This mechanism introduces a learnable gating parameter, $\alpha$, which dynamically modulates the network’s effective depth during training, thereby stabilizing optimization and enabling convergence for deeper architectures. Inspired by this approach, we introduce Residual-Gated Adaptive Kolmogorov–Arnold Networks (RGA KANs), the architecture of which is illustrated in Figure 5.
@@
This basis function design is inspired by the ActLayer [56] and plays a similar role to RFF embeddings used in the PirateNet architecture. In preliminary experiments, RFF embeddings were found to degrade performance in our setting, whereas sine-based KAN layers preserved the benefits of trigonometric features, which have been shown to be particularly effective in many PDE problems [56]. In addition to the trainable coefficients $b^{s}_{jim}$ (initialized using the Glorot-like scheme proposed in this work), we also introduce trainable phase parameters $p_{m}$ (initialized at zero) and frequency parameters $\omega_{m}$ (initially sampled from a standard normal distribution), as in ActLayers.
At this stage, drawing inspiration from the Modified MLP architecture [20], we define two gates using Chebyshev-based KAN layers:
@@
where $\mathbf{y}_{0}$ contains the initial-condition target values and $\mathbf{\tilde{B}}$ is the matrix of basis activations evaluated on the outputs of the final RGA block when the original inputs are set to the collocation points enforcing the initial condition. The vector $\tilde{\mathbf{b}}^{o}$ contains the optimal coefficients obtained by the least-squares fit and can be re-indexed back to the original $(i,m)$ indexing to yield the coefficients $b^{o}_{jim}$ of Eq. (51). Note that, although here we focus on standard PDE benchmarks with initial conditions, this formulation can in principle incorporate arbitrary external data through $\mathbf{y}_{0}$, such as experimental measurements [23].
If non-periodic boundary conditions are present, they can be directly enforced after this final physics-informed layer by multiplying the network output by suitable boundary-shaping functions, thereby ensuring that the solution satisfies these constraints exactly [63]. Based on the above architectural components, for an RGA KAN model with $d_{\text{O}}$-dimensional output, the total number of trainable parameters can be explicitly determined as follows:
@@
Figure 6: Relative $L^{2}$ error across increasing network depths for the Allen–Cahn equation, comparing RGA KANs and cPIKANs. Each column corresponds to a different network width (8, 16, 32) and the number of hidden layers for RGA KANs equals twice the number of RGA blocks. All results are averaged over five random seeds using the proposed Glorot-like initialization scheme. Solid lines indicate mean values and shaded areas denote standard error.
The results demonstrate a significant improvement in stability and accuracy compared to the baseline cPIKANs. Across all widths and depths considered, RGA KANs maintain low relative $L^{2}$ errors without exhibiting divergence, even for the deepest networks. Moreover, a favorable scaling effect is observed: as the network width increases, the relative error decreases consistently, with the widest configuration (width $32$) yielding the best results across depths. More importantly, within each width setting, increasing the number of RGA blocks either preserves performance (plateau behavior) or further reduces the error – a behavior that mirrors the improvements reported for PirateNets over regular MLPs [23].

### v2

@@ 3.2 Small-Scale Benchmarks
To evaluate the effectiveness of the proposed initialization scheme, we compare it against the standard practice for Chebyshev-based KANs. To this end, we conduct experiments on two small-scale benchmarks: (i) function fitting, and (ii) forward PDE problems. For the remainder of this work, we refer to the standard initialization strategy defined in Eq. (18) as the “default” initialization.
@@ 3.2.1 Function Fitting
We first assess initialization performance on five function-fitting tasks of increasing dimensionality: (i) a one-dimensional oscillatory function, $f_{1}(x)$, (ii) the two-dimensional product function, $f_{2}(x_{1},x_{2})$, (iii) a more challenging two-dimensional function borrowed from [MOSS], $f_{3}(x_{1},x_{2})$, (iv) the three-dimensional Hartmann function, $f_{4}(x_{1},x_{2},x_{3})$, which is a common benchmark in function approximation, and (v) the five-dimensional Sobol g-function, $f_{5}(x_{1},\dots,x_{5})$, widely used in global sensitivity analysis. The analytic definitions of all functions are provided in D.
We train Chebyshev-based KANs with polynomial order $D=8$ using both the default and the proposed Glorot-like initialization schemes. We conduct a sweep over architectures of varying width (hidden layer dimensions of 2, 4, 8, 16, 32, 64) and depth (2 to 5 hidden layers) and evaluate performance in terms of the relative $L^{2}$ error with respect to the reference solution. To ensure statistical significance, each configuration is repeated with five random seeds. Full implementation details regarding data generation and optimization settings are provided in C.1. The results of these experiments are summarized in Figure 1. Each heatmap corresponds to one benchmark function, with hidden layer dimension on the horizontal axis and number of hidden layers on the vertical axis. The color scale quantifies the relative improvement achieved by the proposed initialization over the default scheme, computed as
@@
Evidently, the impact of the proposed initialization is substantial in most benchmarks. For the two-dimensional functions ($f_{2}$ and $f_{3}$) and the three-dimensional Hartmann function ($f_{4}$), improvements approach 100% across nearly all architectures, indicating that the proposed initialization reduces the final relative $L^{2}$ error by up to several orders of magnitude compared to the default scheme. The one-dimensional oscillatory function ($f_{1}$) also shows clear gains in the majority of cases, with only two isolated configurations where the default (marginally) outperforms. The five-dimensional Sobol g-function ($f_{5}$) exhibits improvements as well, though they are less pronounced, typically in the range of 5–50%; in this setting, both initializations yield comparable overall accuracy, and therefore the difference in initialization impact is less striking.
Figure 2: Loss throughout training for two representative architectures (top row: three 4-dimensional hidden layers; bottom row: five 16-dimensional hidden layers) across the five benchmark functions. Each subplot shows the mean training loss over five independent runs (solid lines) together with the SEM (shaded area). The final column, corresponding to $f_{5}$, is shown without logarithmic scaling on the $y$-axis, since the loss did not exhibit significant improvement during training.
Apart from the final error metrics, it is also informative to examine the training loss evolution, in order to assess whether the proposed initialization leads to a more effective optimization of the loss function. To this end, we consider two representative architectures: a smaller network with width $4$ and depth $3$, and a larger network with width $16$ and depth $5$. Figure 2 depicts the training loss curves for each of the five benchmark functions under both initialization schemes. Solid lines indicate the mean loss across five independent runs, while the shaded regions correspond to the standard error of the mean (SEM). The proposed initialization consistently accelerates convergence and achieves substantially lower training losses. For the oscillatory function ($f_{1}$) and the two-dimensional cases ($f_{2}$ and $f_{3}$), the difference spans more than two orders of magnitude, with low variability across seeds. Similar improvements are observed for the Hartmann function ($f_{4}$), where the default initialization stalls at higher loss values compared to the proposed scheme. For the Sobol $g$-function ($f_{5}$), both schemes exhibit nearly identical behavior, in line with the earlier observation that this benchmark is less sensitive to initialization. Importantly, these trends are visible in both architectures, demonstrating robustness across different model scales. To provide deeper insight into the stability improvements of the proposed initialization, we also present a Neural Tangent Kernel (NTK) analysis for these architectures in F.1.
@@ 3.2.2 Forward PDE Problems
We next assess the proposed initialization on forward PDE benchmarks using cPIKANs. Specifically, we consider Burgers’ equation as well as the Helmholtz equation with $a_{1}=1$, $a_{2}=4$ (see E for details), following the PIML framework introduced in Section 2, without incorporating any additional adaptive techniques. We use Chebyshev-based KANs of order $D=8$ and conduct a sweep over the same architectural configurations (widths and depths) as in the function-fitting experiments, repeating each experiment with five random seeds. Full details on collocation point distribution and optimization settings are provided in C.1. The results of these experiments are summarized in Figure 3, which combines heatmaps of final relative $L^{2}$ error improvements with representative training-loss curves.
Figure 3: Comparison of default and proposed initialization schemes on Burgers’ (top row) and Helmholtz (bottom row) equations. Left column: heatmaps of relative improvement in final $L^{2}$ error. Middle/Right column: training-loss curves per initialization scheme for a representative architecture of depth = 3/5 and width = 4/16, respectively. Shaded regions denote the SEM across five runs.
As in the function-fitting benchmarks, the proposed initialization consistently outperforms the default scheme in terms of final error, with the exception of only two small architectures where the default initialization yields a marginal advantage. The overall gains are somewhat less pronounced than in Section 3.2.1, which can be attributed to the relatively low number of training iterations and the absence of adaptive training techniques. Nevertheless, the training-loss curves reveal a striking effect of initialization. In particular, for the Helmholtz equation, where it is well known that fast convergence usually requires adaptive weighting of PDE and boundary condition terms (see, e.g., [FAIR]), the proposed initialization achieves up to eight orders of magnitude lower training loss compared to the default scheme. This distinct gap highlights a critical stability advantage: while the default initialization leads to training divergence, the proposed scheme ensures stable convergence toward the solution. This empirical finding is further substantiated by the NTK analysis in F.2, which links the proposed initialization to more favorable spectral properties.
@@ 3.3 Training Divergence with Increasing Depth
The benchmarks considered so far involved relatively shallow networks, consistent with most current KAN applications, where architectures are typically limited to a small number of hidden layers. Within this setting, the proposed initialization was shown to improve accuracy and training stability across both function-fitting and forward PDE tasks. A natural question, however, is whether these gains extend to deeper architectures. To this end, we revisit Burgers’ equation and additionally consider the Allen–Cahn equation (see E for details), training cPIKANs of increasing depth to examine how network depth influences training stability under both initialization schemes.
In contrast to the previous benchmarks, here we adopt a standardized adaptive training pipeline, the configuration and hyperparameter settings of which are provided in C.2, since both PDEs are trained for a large number of iterations and the Allen–Cahn equation in particular fails to yield accurate results without adaptive strategies [Causal, RBA]. We consider Chebyshev-based KANs with polynomial order $D=5$, initialized under both the default and the proposed scheme, with architectures of varying widths $\{8,16,32\}$ and depths ranging from 2 to 12 hidden layers. Performance is evaluated in terms of the relative $L^{2}$ error with respect to the reference solution, with each experiment repeated five times using different random seeds. The results of these experiments are summarized in Figure 4, which presents the relative $L^{2}$ error of each network configuration under both initialization schemes for the two studied PDEs. Each column corresponds to a specific network width, with the horizontal axis of each subplot representing the network depth. Solid lines denote the mean relative $L^{2}$ error over the five independent runs, and the shaded regions indicate the corresponding SEM.
@@
Figure 4: Relative $L^{2}$ error across increasing network depths for Burgers’ (top row) and Allen–Cahn (bottom row) equations, under both default and proposed initialization schemes. Each column corresponds to a different network width (8, 16, 32). Solid lines show mean values over five random seeds, while shaded areas represent the SEM.
The results highlight the consistent advantage of the proposed initialization across all architectures and both PDEs, often yielding relative $L^{2}$ errors lower by several orders of magnitude compared to the default scheme. For the Burgers’ equation, the effect is particularly striking: beyond a depth of four hidden layers, models initialized with the default scheme exhibit complete divergence, with relative errors of $\mathcal{O}(1)$, whereas the proposed initialization maintains stable training and achieves errors as low as $\mathcal{O}(10^{-3})$$\mathcal{O}(10^{-4})$. Although some degradation in accuracy is observed at larger depths for the widest networks (width $32$), no divergence occurs, and the relative errors remain well below those of the default scheme. In contrast, for the Allen–Cahn equation, training instability persists for deeper networks under both initialization schemes. While the proposed initialization consistently improves performance compared to the default, the relative error grows rapidly with depth and reaches $\mathcal{O}(1)$ for the deepest configurations. The lowest errors are typically achieved at shallow depths (i.e., depth $2$ for width $32$ and depth $4$ for the other two cases), after which performance degrades, albeit not monotonically for all cases. Overall, while the proposed scheme substantially improves training stability and accuracy, it does not, by itself, guarantee stable convergence in all cases as network depth increases.
@@ 4 Residual-Gated Adaptive KANs
While the proposed Glorot-like initialization improved training stability and accuracy across all benchmarks, its effectiveness in mitigating depth-related issues proved case-dependent. In particular, for Burgers’ equation, cPIKANs initialized with the proposed scheme remained stable and convergent even at larger depths, whereas for the Allen–Cahn equation similar architectures exhibited divergence beyond a certain number of hidden layers. A similar phenomenon is observed in PINNs: when networks are initialized following the Glorot scheme, training tends to diverge as depth increases, regardless of the specific activation function among those commonly used in practice [PirateNets]. However, before attempting to transfer to cPIKANs the remedies that have been proposed for this behavior in PINNs, it is first necessary to determine whether the underlying mechanisms are indeed analogous.
@@ 4.1 KAN Derivatives at Initialization
In the case of MLP-based networks – the backbone of PINNs – [PirateNets] demonstrated that training divergence with increasing depth originates from the behavior of the network’s derivatives at initialization. To illustrate this, they considered a simplified MLP with scalar input and output, employing the hyperbolic tangent activation $\sigma(x)=\tanh(x)$, and focused on the first-order derivative with respect to the input. At initialization, the network operates in a near-linear regime where $\sigma(x)\approx x$, leading to the following approximation for the first-order derivative of Eq. (10) with respect to the input coordinate $x$:
@@
For Chebyshev-based KANs, where the basis functions are given by Eq. (16), it can be shown (see B.2 for the detailed derivation) that, in the linear regime,
@@ 4.2 Proposed Architecture
To address these instabilities in PINNs, [PirateNets] introduced the PirateNet architecture (see G for details). PirateNets incorporate several architectural components known to improve accuracy, such as random Fourier feature (RFF) embeddings [RFF] and a physics-informed initialization of the final network layer. However, the key idea for resolving the depth-related issue is the introduction of an adaptive skip connection. This mechanism introduces a learnable gating parameter, $\alpha$, which dynamically modulates the network’s effective depth during training, thereby stabilizing optimization and enabling convergence for deeper architectures. Inspired by this approach, we introduce Residual-Gated Adaptive Kolmogorov–Arnold Networks (RGA KANs), the architecture of which is illustrated in Figure 5.
@@
This basis function design is inspired by the ActLayer [ActNet] and plays a similar role to RFF embeddings used in the PirateNet architecture. In preliminary experiments, RFF embeddings were found to degrade performance in our setting, whereas sine-based KAN layers preserved the benefits of trigonometric features, which have been shown to be particularly effective in many PDE problems [ActNet]. In addition to the trainable coefficients $b^{s}_{jim}$ (initialized using the Glorot-like scheme proposed in this work), we also introduce trainable phase parameters $p_{m}$ (initialized at zero) and frequency parameters $\omega_{m}$ (initially sampled from a standard normal distribution), as in ActLayers.
At this stage, drawing inspiration from the Modified MLP architecture [Wang1], we define two gates using Chebyshev-based KAN layers:
@@
where $\mathbf{y}_{0}$ contains the initial-condition target values and $\mathbf{\tilde{B}}$ is the matrix of basis activations evaluated on the outputs of the final RGA block when the original inputs are set to the collocation points enforcing the initial condition. The vector $\tilde{\mathbf{b}}^{o}$ contains the optimal coefficients obtained by the least-squares fit and can be re-indexed back to the original $(i,m)$ indexing to yield the coefficients $b^{o}_{jim}$ of Eq. (44). Note that, although here we focus on standard PDE benchmarks with initial conditions, this formulation can in principle incorporate arbitrary external data through $\mathbf{y}_{0}$, such as experimental measurements [PirateNets].
If non-periodic boundary conditions are present, they can be directly enforced after this final physics-informed layer by multiplying the network output by suitable boundary-shaping functions, thereby ensuring that the solution satisfies these constraints exactly [ExactBCs]. Based on the above architectural components, for an RGA KAN model with $d_{\text{O}}$-dimensional output, the total number of trainable parameters can be explicitly determined as follows:
@@
Figure 6: Relative $L^{2}$ error across increasing network depths for the Allen–Cahn equation, comparing RGA KANs and cPIKANs. Each column corresponds to a different network width (8, 16, 32) and the number of hidden layers for RGA KANs equals twice the number of RGA blocks. All results are averaged over five random seeds using the proposed Glorot-like initialization scheme. Solid lines indicate mean values and shaded areas denote the SEM.
The results demonstrate a significant improvement in stability and accuracy compared to the baseline cPIKANs. Across all widths and depths considered, RGA KANs maintain low relative $L^{2}$ errors without exhibiting divergence, even for the deepest networks. Moreover, a favorable scaling effect is observed: as the network width increases, the relative error decreases consistently, with the widest configuration (width $32$) yielding the best results across depths. More importantly, within each width setting, increasing the number of RGA blocks either preserves performance (plateau behavior) or further reduces the error – a behavior that mirrors the improvements reported for PirateNets over regular MLPs [PirateNets].